OpenStudy (anonymous):

An equation of a circle is x2 + y2 + 10x – 6y + 18 = 0. Show all your work in determining the center and radius of this circle. In complete sentences, explain the procedure used.

OpenStudy (apoorvk):

Hey complete squares in 'x' and 'y'.

OpenStudy (anonymous):

I dont know how?

OpenStudy (apoorvk):

Now, when you complete squares, you will get the equation in the form of: \[(x-a)^2 + (y-b)^2 = r^2\] Here, (a,b) would be the center and 'r' the center.

OpenStudy (anonymous):

ok? but i dont know how to do that!

terenzreignz (terenzreignz):

First, put the x's and y's together: \[x^{2} +10x + y^{2} - 6y +18 = 0\] I'll show you how to complete the square for the 'x' part, you do the same for the 'y' Let's only look at the x's for now: \[x^{2} + 10x\] What we do is we take the term with just the x (not x squared) and look at its coefficient, which happens to be 10 (whether or positive or negative is unimportant, at this point) Now pay attention to this: We take that coefficient (10), we divide it by two (5) and then square it (25) Now that we have that, we ADD it to this expression to get \[x^{2} + 10x + 25\] BUT, we have changed the expression by adding 25, to make sure it's the same expression, we also subtract 25, like so: \[x^2 + 10x + 25 -25\] Look at the first three terms.... As you can see, it is a PERFECT SQUARE TRINOMIAL, so this can be written as \[(x^2 + 10x + 25) -25\] \[(x + 5)^{2} - 25\] And that's how you complete the square, now try doing that with the 'y' part :D